# Black Big Boxes: Do Language Models Hide a Theory of Adjective Order?

### v1

@@ Conclusion
The Pythia LMs acquire a strong sense of adjective order that is learned early in training at around 1.5% of the total amount of training. We showed that sentence context has a positive effect on AOP, bumping up AOP-from 91% to around 94%. By localizing the AOP scores on the token-level we show how these order preferences manifest at different points in a phrase. In the next experiments we will investigate these results in more detail, connecting LM AOP to cognitive predictors and training data statistics.
5 Experiment 2: Cognitive Predictors
@@
Metric AOP-$\rho(\textup{AOP-}\Delta(\bullet))$ $\rho(\textup{AOP-}\Delta(\bullet|c))$
AOP-$\Delta(\bullet)$ 91.6 0.71
@@
[…]
PMI(a;n) 62.4 0.27 0.20
Subj(a) 69.3 0.07 0.08
[…]
%$\%(\textsc{a}_{1}\textsc{a}_{2})$ 90.3 0.58 0.53
%$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ 87.6 0.54 0.53
Table 1: Predictive accuracy of adjective order on CAP for various metrics: AOP-$\Delta$ for Pythia-12b, 3 cognitive predictors (§5), and relative n-gram counts from The Pile (§6). The highest accuracy for each group is bolded. For each metric we provide the Spearman correlations with respect to the AOP-$\Delta$ scores of Pythia-12b.
As discussed in §2.1, a range of cognitively plausible predictors have been proposed to explain adjective order. In the previous section we have shown that LMs are highly capable of predicting adjective order, and in this section we investigate how their predictions are related to various theories on adjective order. We evaluate the performance of three cognitive predictors—Length, PMI, and Subjectivity— and correlate them against AOP-$\Delta$ scores across training.
5.1 Cognitive Predictors
For each predictor we compute its score for a particular adjective pair. An adjective order predictor $f$ returns a graded score for adjective order, akin to our formulation of AOP-$\Delta$.
Length
For Length—which hypothesizes that the first adjective is the shortest—the score for adjectives a1 and a2 is expressed as
$f(\textsc{a}_{1},\textsc{a}_{2})=|\textsc{a}_{2}|-|\textsc{a}_{1}|$
We also considered expressing length as the number of tokenizer subwords, since this is the only notion of length LMs have access to, but found that this led to similar results as string length.
PMI
We follow the procedure of Futrell et al. (2020a) for computing PMI, which is based on the distribution over adjective relations in dependency parse trees. It is defined as
$\textup{PMI}(\textsc{a};\textsc{n})=P(\textsc{a},\textsc{n})/(P(\textsc{a})\cdot P(\textsc{n}))$
$P(\textsc{a},\textsc{n})$ is computed as the normalized number of occurrences of adjective a and noun n in amod relation. $P(\textsc{a})$ and $P(\textsc{n})$ are obtained by marginalizing over this joint distribution. PMI hypothesizes that adjectives closer to the noun have a higher PMI:
$f(\textsc{a}_{1},\textsc{a}_{2},\textsc{n})=\textup{PMI}(\textsc{a}_{2};\textsc{n})-\textup{PMI}(\textsc{a}_{1};\textsc{n})$
Subjectivity
We leverage the 450 subjectivity ratings provided by Dyer et al. (2023), to compute subjectivity scores for the subset of adjective pairs that contain a rating. This procedure retains 37.5% of the adjective pairs in CAP (1005 items). Subjectivity as predictor hypothesizes that more subjective adjectives occur first (Scontras et al., 2017):
$f(\textsc{a}_{1},\textsc{a}_{2})=\textup{Subj}(\textsc{a}_{1})-\textup{Subj}(\textsc{a}_{2})$
5.2 Predictive Accuracy
We present the predictive accuracy of each predictor in Table 1. Subjectivity is the strongest predictor with 69.3% accuracy, followed by PMI (62.4%) and Length (49.6%). While Subjectivity reaches the highest accuracy, its correlation with the LM’s AOP-$\Delta$ scores is surprisingly the lowest of the three. We plot these correlations over time in Appendix A.2: all three correlations develop in the acquisition phase and remain stable in the consolidation phase.
We also compute the overlap between these three predictors on the 1005 items for which we have subjectivity scores:
[Uncaptioned image][Uncaptioned image]
The first Venn diagram makes clear that the cognitive predictors have low agreement: only 213 orders are predicted correctly by all three. The broad coverage (93.1%) of the union, however, shows that most of the CAP items are explained by at least one of these predictors. In the four diagrams on the right we show the overlap of the predictors with the AOP-$\Delta$ scores of Pythia-12b. All three predictors and their intersection are almost fully subsumed by the AOP-$\Delta$ predictions. This demonstrates that AOP-$\Delta$’s unique contribution goes significantly above that of the ensemble of the cognitive factors, raising the question whether LMs ‘hide’ a meaningful linguistic theory that extends beyond established factors.
6 Experiment 3: Is AOP Driven by Data Statistics?
@@
In order to truly understand what LMs base their adjective order preferences on, we need to get a better insight into how double adjective constructions are distributed in the training data. The way that contemporary LMs are trained nowadays makes this challenging: many models are released without open access to their training data, and the scale of training corpora that are available poses a challenge for targeted analysis.
6.1 WIMBD
To overcome this, we make use of the WIMBD API (What’s In My Big Data, Elazar et al., 2023), which provides a fully indexed search engine over various large-scale corpora such as The Pile, on which the Pythia models have been trained. WIMBD makes it possible to retrieve the exact $n$-gram statistics that LMs have been trained on, and as such provides a detailed insight into the collocational factors that play a direct role in shaping AOP.
%$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})=\log\#(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-\log\#(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$
with similar formulations for adjective bigrams (%($\textsc{a}_{1}\textsc{a}_{2}$)) and unigrams (%($\textsc{a}_{1}$)).
Predictive Accuracy
We present the predictive accuracy of these relative n-gram occurrences in Table 1. %(a1) predicts adjective order with 60.5%, which provides weak evidence that more frequent adjectives tend to occur first (Martin, 1969; Trotzke and Wittenberg, 2019). %(a1a2) is the strongest predictor of adjective order—except for the LMs—at 90.3%. This shows that a large majority of the items in CAP can be resolved based on bigram statistics alone, which is leveraged by the LMs during training. However, n-gram collocations alone do not fully explain LM performance: a large amount of trigrams are not part of the Pile and performance for %$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ drops to 87.6%.
@@ 6.1 WIMBD
To investigate how n-gram statistics play a role in the formation of AOP-$\Delta$ during training, we compute the Spearman correlation between various log-transformed n-gram counts and a LM’s AOP-$\Delta$ on intermediate checkpoints. We plot this in Figure 4, for contextual and isolated AOP-$\Delta$. The trends we observe here align closely with the three learning phases that we outlined in §4. The initialization phase (1) is marked by a rapid acquisition of unigram and bigram statistics, that remain unaffected by context at this point. Then, at the start of the acquisition phase (2), in which AOP performance rapidly improves, the model starts to incorporate linguistic context in its uni- and bigram predictions. Finally, in the consolidation phase (3), trigram predictions start being affected by context as well, whereas unigram and bigram predictions remain relatively stable. The fully trained model strongly reflects these correlations to data statistics, which explains the strong relation we observed in Figure 1 between bigram frequencies and AOP-$\Delta$.
@@ 6.2 Generalizing to Unseen Adjective Orders
Figure 5: The contextual AOP-performance for Pythia-70m and 1.4b across training, split out for items that have been seen 0, 1, 2 to 10, and more than 10 times at each specific checkpoint.
Having access to the n-gram distribution a model was trained on allows us to investigate its performance on unseen adjective orders: is the model able to generalize from an abstract signal of adjective order to novel orders? However, due to the enormous size of The Pile (300B tokens) it turns out that number of unseen adjective pairs in CAP is too small to draw significant conclusions from (only 4 out 2678 pairs). This leaves us with two options: either we collect an additional sample of rare adjective pairs that are not part of the Pile, or we investigate the model at intermediate checkpoints where a greater amount of adjective pairs has not been encountered yet. An issue with collecting highly rare pairs is that their canonical order will be much harder to determine, for which we would require extensive human judgments on the most natural order. We therefore go for the second option, and set up a procedure to measure the data statistics during training.
@@
Since WIMBD only provides $n$-gram statistics for The Pile as a whole, we need to collect the intermediate statistics ourselves. Using the Pythia batch viewer (Biderman et al., 2023b) this becomes possible: our implementation is able to process the first 10% of The Pile (up to batch 14,000; 29.4B tokens) in around 3 hours. We collect the bi-gram counts of all adjective pairs in CAP on the batch level, allowing us to determine exactly when and how often a specific adjective pair has been seen by the model.
@@
Based on these intermediate $n$-gram counts, we compute the AOP-scores for various splits based on the number of times an adjective pair has been encountered. We focus on the cases where the swapped order has never been encountered: this allows us to isolate the moment a model encounters an adjective combination for the first time, and measure its ‘zero-shot’ performance on cases that are not already skewed towards to swapped order. For each model checkpoint we collect the $n$-gram counts at that particular point in training, and split the CAP items based on the number of times an a1a2 bi-gram has been seen. We provide the size of these splits across training in Appendix A.3.
We present the results for Pythia-70m and 1.4b in Figure 5. The performance on unseen pairs is remarkably high, reaching around 85% at the highest point for both models. Interestingly, the gap between unseen pairs and pairs seen once is far greater for the 1.4b model than the small 70m model. This shows that this larger model is able to acquire adjective order preference faster: a single exposition to an order immediately leads to an increase in AOP-$\Delta$. We provide a more detailed item-level analysis for these items in Appendix A.4: while it is clear on the split level that a single occurrence improves AOP-%, we do not observe such a clear effect on the item level. Interestingly, as the number of unseen adjective pairs decreases, the performance on unseen pairs decreases as well. This shows that the predicting the order for the set of unseen pairs around batch 2,000 (6% of CAP) could be better inferred from a general signal of adjective order than the set at batch 14,000 (1.3% of CAP).
@@ Conclusions
The question we set out to answer in this set of experiments is whether AOPs are driven by data statistics. We approached this question both in relation to the $n$-gram distribution of the Pile and to the shifting $n$-gram distribution across training. On the one hand, AOP is strongly driven by data statistics: we find a strong correlation between AOP-$\Delta$ and count-based AOP scores, which is also visualized in Figure 1. However, the relatively strong performance on the set of unseen adjective pairs during training indicates that LMs are able to infer order from a more general signal for adjective pairs it has not encountered before.
7 Experiment 4: The Role of Context
Figure 6: A: Item-level AOP-$\Delta$ scores with and without sentence-context for Pythia-12b. We provide a sample of sentences from each quadrant in Appendix A.1. B: The relative impact ($c$-$\Delta$, Eq. 5) of an item’s true context compared to an expectation over random contexts (Eq. 4).
@@ 7 Experiment 4: The Role of Context
Contextual Improvements
To gain a better insight how much items are impacted by context, we investigate AOP-$\Delta$ scores on an item-level with and without context. In Figure 6A, we plot the AOP-$\Delta$ scores in isolation against the contextual scores, and mark the four quadrants of items that result in a positive or negative score after adding context. In the majority of cases (89.4%), AOP-$\Delta$ is positive both with and without context: the model is able to determine adjective order based on the adjective-noun triplet alone. In 4.8% of the cases, however, adding context flips a negative AOP-$\Delta$ score to a positive one. Adding context has a boosting effect in general: in 65.1% of the cases, AOP-$\Delta$ is increased by adding context. We provide an overview of how these ratios are affected by model size in Appendix A.5: the number of cases where context improves AOP-is actually larger for small models, but so is the number of cases where context worsens it.

### v2

@@ Conclusion
The Pythia LMs acquire a strong sense of adjective order that is learned early in training at around 1.5% of the total amount of training. We showed that sentence context has a positive effect on AOP, bumping up AOP-% from 88% to around 91%. By localizing the AOP scores on the token-level we show how these order preferences manifest at different points in a phrase. In the next experiments we will investigate these results in more detail, connecting LM AOPs to training data statistics.
5 The Role of Training Data Statistics in AOP
Having established that language models achieve high accuracy on adjective order prediction, we now examine the extent to which this behaviour can be explained by distributional properties of the training data. Specifically, we ask to what extent simple co-occurrence statistics explain model predictions, and where systematic deviations from such statistics arise.
To ensure that our subsequent analyses are not specific to a single model family, we first examine adjective order behaviour across a range of contemporary large language models. Table  5 shows the performance of four contemporary, mid-size, open-source LLMs: Llama3-8b, OLMo2-7b, OLMo3-7b and Qwen3-8b (Grattafiori et al., 2024; Yang et al., 2025; OLMo et al., 2025a, b). These models perform similarly to Pythia (slightly worse, at 89 to 91%) with high AOP-$\Delta$ correlation, supporting the use of the Pythia models as a representative case in our subsequent analyses.
As a next step, we assess whether LM adjective order behaviour can be accounted for by cognitive predictors proposed in prior work (§2.1). Compared to such predictors, language models exhibit adjective order behaviour that is not well explained by any single factor. In Table 5 we show the accuracy of two predictors: PMI between adjective and noun (62.4%), and adjective subjectivity (69.3%). As show in Appendix A.2, these predictors show limited overlap with the AOP-$\Delta$ scores, indicating that LM behaviour cannot be straightforwardly reduced to the effects captured by these individual predictors.
@@ 5 The Role of Training Data Statistics in AOP
Table 1: Predictive accuracy of adjective order on CAP for various metrics: AOP-$\Delta$ for Pythia-12b and relative $n$-gram counts from The Pile (§5). The highest accuracy for each group is bolded. For each metric we provide the Spearman correlations with respect to the AOP-$\Delta$ scores of Pythia-12b. AOP scores for the four LLMs denote contextual AOP.
5.1 Infini-gram
In order to better understand the factors underlying LM adjective order behaviour, we examine how double adjective constructions are distributed in the training data. The way that LMs are trained nowadays makes this challenging: many models are released without open access to their training data, and the scale of training corpora that are available poses a challenge for targeted analysis.
To overcome this, we make use of the Infini-gram API (Liu et al., 2024), which provides a fully indexed search engine over various large-scale corpora such as The Pile, on which the Pythia models have been trained. Infini-gram makes it possible to retrieve the exact $n$-gram statistics that LMs have been trained on, and as such provides a detailed insight into the collocational factors that play a direct role in shaping AOP.
$\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})=\log\#(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})-\log\#(\textsc{a}_{2}\textsc{a}_{1}\textsc{n})$
@@ 5.1 Infini-gram
Predictive Accuracy
We present the predictive accuracy of these relative $n$-gram occurrences in Table 5. %(a1) predicts adjective order with 55.2%, which provides weak evidence that more frequent adjectives tend to occur first (Martin, 1969; Trotzke and Wittenberg, 2019). %(a1a2) is the strongest predictor of adjective order at 86.8%. This shows that a large majority of the items in CAP can be resolved based on bigram statistics alone, which is leveraged by the LMs during training. However, $n$-gram collocations alone do not fully explain LM performance: a large amount of trigrams are not part of the Pile and performance for $\%(\textsc{a}_{1}\textsc{a}_{2}\textsc{n})$ drops to 76.7%.
@@
To investigate how $n$-gram statistics relate to the formation of AOP-$\Delta$ during training, we compute the Spearman correlation between various log-transformed $n$-gram counts and a LM’s AOP-$\Delta$ on intermediate checkpoints. We plot this in Figure 4, for contextual and isolated AOP-$\Delta$. The trends we observe here align closely with the three learning phases that we outlined in §4. The initialization phase (1) is marked by a rapid acquisition of unigram and bigram statistics, that remain unaffected by context at this point. Then, at the start of the acquisition phase (2), in which AOP performance rapidly improves, the model starts to incorporate linguistic context in its uni- and bigram predictions. Finally, in the consolidation phase (3), trigram predictions start being affected by context as well, whereas unigram and bigram predictions remain stable. The fully trained model strongly reflects these correlations to data statistics, which explains the strong relation we observed in Figure 1 between bigram frequencies and AOP-$\Delta$. This successive alignment from unigram to higher $n$-gram statistics has also been observed in prior work (Chang and Bergen, 2022; Michaelov et al., 2025).
@@ 5.2 Generalizing to Unseen Adjective Orders
Having access to the $n$-gram distribution a model was trained on allows us to investigate its behaviour on adjective orders not previously observed during training. We ask whether models exhibit systematic generalization to unseen adjective combinations, beyond direct memorization of observed co-occurrence patterns. However, due to the enormous size of The Pile (300B tokens) it turns out that number of unseen adjective pairs in CAP is too small to draw significant conclusions from (only 39 out 9396 pairs, 0.4%). This leaves us with two options: either we collect an additional sample of rare adjective pairs that are not part of the Pile, or we investigate the model at intermediate checkpoints where a greater amount of adjective pairs has not been encountered yet. An issue with collecting highly rare pairs is that their canonical order will be much harder to determine, for which we would require extensive human judgments on the most natural order. We therefore go for the second option, and set up a procedure to measure the data statistics during training.
Figure 5: The contextual AOP-% performance for Pythia-70m and 1.4b across training, split out for items that have been seen 0, 1, 2 to 10, and more than 10 times at each specific checkpoint. We provide the size of these splits across training in Appendix A.3.
@@
Since Infini-gram only provides $n$-gram statistics for The Pile as a whole, we need to collect the intermediate statistics ourselves. Using the Pythia batch viewer (Biderman et al., 2023b) this becomes possible: our implementation is able to process the first 10% of The Pile (up to batch 14,000; 29.4B tokens) under one hour on a consumer-grade laptop. We collect the bi-gram counts of all adjective pairs in CAP on the batch level, allowing us to determine exactly when and how often a specific adjective pair has been observed during training.
@@
Based on these intermediate $n$-gram counts, we compute the AOP-% scores for various splits based on the number of times an adjective pair has been encountered. We focus on the cases where the swapped order has never been encountered: this allows us to isolate the moment a model encounters an adjective combination for the first time, and measure its ‘zero-shot’ performance on these cases. For each model checkpoint we collect the $n$-gram counts at that particular point in training, and split the CAP items based on the number of times an a1a2 bi-gram has been seen.
We present the results for Pythia-70m and 1.4b in Figure 5. The performance on unseen pairs is remarkably high, reaching around 85% at the highest point. We provide a more detailed item-level analysis for these items in Appendix A.4: while it is clear on the split level that a single occurrence improves AOP-%, we do not observe such a clear effect on the item level. Interestingly, as the number of unseen adjective pairs decreases, the performance on unseen pairs decreases as well. As training progresses and the pool of unseen adjective pairs becomes smaller, performance on these unseen items decreases. Since the set of unseen pairs at later checkpoints is a strict subset of earlier unseen sets, this trend likely reflects increasing item difficulty rather than a qualitative change in model generalization.
@@ Conclusions
The question we set out to address in this set of experiments is the extent to which adjective order behaviour in language models can be explained by direct co-occurrence statistics. We approached this question by analysing the overall $n$-gram distribution of the Pile and its evolution over the course of training. On the one hand, AOP is strongly connected to training data statistics: we observe a strong correlation between AOP-$\Delta$ and count-based AOP scores, as visualized in Figure 1. On the other hand, the relatively strong performance on adjective pairs not previously observed during training indicates that models generalize adjective order behaviour beyond direct memorization of seen co-occurrences.
Figure 6: A: Item-level AOP-$\Delta$ scores with and without sentence-context for Pythia-12b. We provide a sample of sentences from each quadrant in Appendix A.1. B: The contextual gain ($c$-$\Delta$, Eq. 5) of an item’s true context compared to an expectation over random contexts (Eq. 4).
6 The Role of Context
@@
6.1 Quantifying Context Impact
Context improves predictions a lot
To understand in what ways AOP is impacted by context, we investigate AOP-$\Delta$ scores on an item-level with and without context. In Figure 6A, we plot the AOP-$\Delta$ scores in isolation against the contextual scores, and mark the four quadrants of items that result in a positive or negative score after adding context. In the majority of cases (85.2%), AOP-$\Delta$ is positive both with and without context: the model is able to determine adjective order based on the adjective-noun triplet alone. In 6.5% of the cases, however, adding context flips a negative AOP-$\Delta$ score to a positive one. Adding context has a boosting effect in general: in 67.6% of the cases, AOP-$\Delta$ is increased by adding context. We provide an overview of how these ratios are affected by model size in Appendix A.5: the number of cases where context improves AOP-% is actually larger for small models, but so is the number of cases where context worsens it.
Figure 7: Contextual AOP-% and Context Gain against Sentence Length and (Adjective) Order Entropy for all Pythia models. Scores are binned based on quantiles: $x$-tick labels indicate the lower bound of each bin (e.g. $[3,6)$ for the first tick).
@@ 6.1 Quantifying Context Impact
$\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}\mid\mathcal{C})=\mathbb{E}_{c\sim\mathcal{C}}\left[\textup{AOP-}\Delta(\textsc{a}_{1}\textsc{a}_{2}\textsc{n}\mid c)\right]$ (4)
We compute the expectation over a random sample of 500 contexts, taken from CAP ($\mathcal{C}$). To express the relative impact of adding context, we regress out a model’s isolated AOP-$\Delta$ score by subtracting it from the contextual scores, which we call the contextual gain ($c\textup{-}\Delta$):
By computing contextual gain for both the original context and for the expectation over random contexts we can quantify how much more important the original context is for AOP-$\Delta$.
We plot the results for this experiment in Figure 6B. The true context improves isolated AOP-$\Delta$ in 65.1% of the cases, whereas random contexts lead to an improvement for only 39.6% (for 33.2% only the true context leads to an improvement whereas random contexts don’t).
Figure 8: Four sentences annotated with the Integrated Gradient attribution scores, that each denote a specific different way in which part-of-speech plays a role in contextual adjective order.
Context Properties
In Figure 7 we plot the impact of context length against AOP-% and Context Gain. Longer contexts lead to better performance as a whole (7A), and lead to larger differences between the contextual and isolated AOP scores (7C). This effect gradually increases with model size: longer contexts have the greatest effect for the largest model.
One way that contextual cues may help a model is to disambiguate adjective orders with high entropy. We compute adjective order entropy based on the $\textsc{a}_{1}\textsc{a}_{2}$ bigram counts from Infini-gram (§5.1). In Figure 7B we demonstrate that model performance drops significantly for adjective pairs with high entropy: these are the hardest cases to get right. Surprisingly, however, the context gain does not increase for high entropy items (7D), indicating that the model can not leverage contextual cues for disambiguation.
6.2 Uncovering Context Cues
Thus far we have only examined the role of context on a macro-level. We now present an approach to identify specific words in the context that are of importance to the model’s predictions. For this we make use of feature attribution methods, which express the relative importance of input features to a prediction.
Procedure
Our procedure for identifying salient context cues is inspired by Sarti et al. (2024), who use feature attribution techniques to identify contextual cues in machine translation. Key to this approach is the usage of contrastive explanations (Yin and Neubig, 2022), where the contribution of a context token is expressed with respect to the difference of two possible outputs (in our case the two adjective orders).
  1. 1.
    Using the context gain scores of Eq. 5, we create a subset of context-dependent constructions by taking the top 20% of items in CAP. These are the items for which the sentence context had the greatest impact on the model’s adjective order preference.
  2. 2.
    We compute the contribution of the tokens in the context using Integrated Gradients (Sundararajan et al., 2017) with respect to the AOP-$\Delta$ scores. [^2] Scores are obtained using inseq (Sarti et al., 2023), with 0-valued baselines. We sum over the embedding dimension to aggregate subword-level attributions, and sum subword attributions to obtain token-level attributions. Since Integrated Gradients is linear we can obtain contrastive explanations by subtracting the attribution scores of the swapped phrase from the scores of the original phrase. Positive contribution scores indicate that that token boosts the original order, whereas a negative score indicates a contribution to the opposite order. [^3] Feature attribution methods have been under extensive scrutiny in recent years: the complexity of large-scale models makes it challenging to capture their behaviour in a singular importance score, which will lead to disagreement between different methods (Neely et al., 2022). Integrated Gradients have nonetheless been shown to be reliable estimators of feature importance (Mariotti et al., 2024; Sarti et al., 2024), and therefore serve as a useful baseline for our experimental setup. We acknowledge though that different attribution methods may yield different insights.
  3. 3.
    We filter for cases where the top contributing token had a score of at least 1, indicating that there is a salient cue present in the context.
  4. 4.
    In order to map our results to a ‘global’ interpretation we POS tag the contexts using SpaCy’s en_core_web_lg tagger (Honnibal et al., 2020). This allows us to express the model behaviour in terms of POS classes, rather than on a lexical level.
Figure 9: Distribution of the most important token in the context, grouped by POS tag. Each point indicates the number of times that POS tag was the most important token, and the median (relative) distance to the adjective phrase. POS tags with a frequency below 30 are filtered out.
Results
We perform our procedure on the Pythia-1.4b model. In Figure 8, we show four annotated sentences that are sampled randomly from the corpus. Each sentence exemplifies the way different part-of-speech classes can play a role in influencing contextual AOP. It can be seen that a context may contain ‘competing’ cues that prefer the natural and swapped adjective order, highlighting that context can impact model behaviour in different ways.
To aggregate this result, we measure the amount of times a POS tag was the most important token in the context, and its distance from the adjective phrase. We visualize this in Figure 9. From this plot we conclude that there are two types of contextual cues that play a role:
  1. 1.